cold storage compartment is 4 m diameter and 4.5 m long with a thickness of 75 mm with insulating material which has a coefficient of thermal conductivity of 5.8 x10^-3 W/m-K. Calculate the quantity of heat energy leaking through the insulation per hour when the outside and inside temperature of the insulation are 15 R and -5 C.


1
Expert's answer
2021-06-07T09:33:26-0400

Gives

Thickness(x)=75mm

Length (h)=4.5m

Radius (r)=2m

Thermal conductivity "K=5.8\times10^{-3}W/m-K"

Area(A)="\pi r^2h"

"A=3.14\times4\times4.5=56.52m^2"

Out side temperature (To)=25R=18.75°C

Inside temperature (Ti)=-5°C

Now we can written as

Rate of heat


"\frac{Q}{t}=\frac{KA(T_o-T_i)}{x}"

Put value

"\frac{Q}{t}=\frac{5.8\times10^{-3}\times{56.52}(18.75-(-5))}{75\times10^{-3}}"

"\frac{Q}{t}=103.809" J/sec

Heat leakage per hour

"Q=\frac{103.809}{3600}=0.0289" J/hour